OpenStudy (anonymous):

Okay, first question: f(g(x)) when f(x)=(2x-2)/(2x-1) and g(x)=1/(4x) the answer should be (4x-1)/(2x-1) but i dont know how. please help

OpenStudy (kkutie7):

\[\frac{2(\frac{1}{4x})-2}{2(\frac{1}{4x})-1}\] plug g(x) in for the x int F(G(x))

OpenStudy (kkutie7):

\[\frac{\frac{2}{4x}-2}{\frac{2}{4x}-1}\] \[\frac{\frac{1}{2x}-2}{\frac{1}{2x}-1}\] \[\frac{\frac{1}{2x}-\frac{4x}{2x}}{\frac{1}{2x}-\frac{2x}{2x}}\] \[\frac{\frac{1-4x}{2x}}{\frac{1-2x}{2x}}\]

OpenStudy (kkutie7):

\[\frac{1-4x}{2x}*\frac{2x}{1-2x}\] \[\frac{2x-8x^{2}}{2x-4x^{2}}\] \[\frac{1-4x}{1-2x}\]